OpenStudy (anonymous):

Just want my answer checked.. is the derivative of x/x^2 +x = -1/(x^2+3)^2 ???

OpenStudy (anonymous):

is not, what should my answer be so i can work through it.

OpenStudy (anonymous):

i've gotten\[\frac{-x^2}{(x^2+x)^2}\] i'll do it again with quotient to double check

OpenStudy (anonymous):

ok thxx/

OpenStudy (anonymous):

yeah mine checks out. If you prefer product rule (i often do) you can rearrange first to avoid the quotient:\[\frac{x}{x^2+x}=(x)*(x^2+x)^{-1}\]

OpenStudy (anonymous):

i re-did it and for the sane asnswer you did as a negative, should it be negative or postitive?

OpenStudy (anonymous):

I got a negative. my second post has a postive but is the original question.

OpenStudy (anonymous):

ooh okay thanks a lot !!